OpenStudy (anonymous):

Which of the following functions decreases, going downwards from left to right on a graph? A. y = 4x B. y = 5x C. y = (1/2)^x D. y = 3(2)^x

OpenStudy (anonymous):

@ninjawolfjj

OpenStudy (anonymous):

I feel like you are missing a negative sign...

OpenStudy (anonymous):

Well...

OpenStudy (anonymous):

hmm let me check

OpenStudy (anonymous):

no not missing anything:)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

would my answer be C?

OpenStudy (anonymous):

I believe it is probably C.

OpenStudy (mathstudent55):

C is correct. Since you are raising a fraction between 0 and 1 to a power, the higher the power, the smaller the number will be. For example: For x = 1, 2, 3 (1/2)^1 = 1/2 (1/2)^2 = 1/4 (1/2)^3 = 1/8 As you can see, as x increased, from 1 to 2 to 3, f(x) decreased from 1/2 to 1/4 to 1/8